Adv. Math. Econ. 5, 65-96 (2003)

Implicit functions and diffeomorphisms without C^{1}

Leonid Hurwicz^{1} and Marcel K. Richter^{2}

1 Department of Economics, University of Minnesota, 271 19th Avenue South,

Minneapolis, MN 55455, USA 2 Department of Economics, University of Minnesota, 271 19th Avenue South,

Minneapolis, MN 55455, USA

Received: August 1, 2002 Revised: August 22, 2002 JEL classification: C600, C650 Mathematical Subject Classification (2000): 26B10, 58C15 Abstract. We prove implicit and inverse function theorems for non-C^{1} functions, and characterize non-C^{1} diffeomorphisms. Key words: Implicit function theorem, inverse function theorem, diffeomorphism, non-C^{1} analysis